# Mutation-Guided Unit Test Generation with a Large Language Model

### v4

@@ Abstract
Index Terms:
Unit Test Generation, Mutation Testing, Large Language Models
@@ I Introduction
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort [48, 30, 22, 10]. With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing [53, 36, 12, 50, 54] and search-based methods [29, 59, 58, 65]—to LLM-based approaches for generating unit tests [27, 66, 37].
@@
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies [18, 11, 39], high code coverage does not necessarily imply strong fault detection capability. Other research [18, 9, 28] suggests that the mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants to the total number of generated mutants.
@@
Recent efforts have begun exploring mutation-based evaluation in the context of LLM-generated test cases. For example, Foster et al. [28] investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases using LLMs based on existing ones, and on generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. [23] targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze the reasons why live mutants remain undetected, nor does it consider uncovered mutants, which are challenges that we address in our study.
@@
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly aims to maximize mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies [63, 68, 55], incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases, caused by incorrect usage of assertion functions, including erroneous invocations and the use of inappropriate parameters.
@@
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Last, since achieving maximum mutation score is inherently difficult, we further introduce an iterative test generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM. Like many recent studies [72, 34, 60], our approach focuses on function-level testing. While this prior work also generates test cases for functions from complex datasets such as Defects4J [41], they handle cross-module dependencies explicitly by just providing them in the prompts. In contrast, our work does not aim to address dependency resolution at this stage, but instead focuses on making significant improvements in mutation scores for function-level testing. Nevertheless, as discussed below, one of our datasets includes solutions to challenging algorithmic problems with complex program structures, providing a more complex dataset for evaluating the effectiveness of MutGen in generating test cases.
@@
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work [63, 23, 61], and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
@@ II Motivating Example and Informal Description
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work [28, 18, 9], high coverage does not necessarily imply strong fault detection capability, for example, when measured as mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 2 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of data inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases such as February 29th on leap years, likely due to the absence of such examples in the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the comments provided within the program under test.
Further, when relying on mutation instead of code coverage, the generated inputs often fail to kill mutants effectively. For example, when the mutation operator changes a conditional boundary and alters line 24 to:
Listing 1: Example mutant introduced at line 24
@@
Listing 2: Example test function generated by Llama-3.3
@@ II-A Preprocessing Stage
In this example, Llama-3.3 returns the following summary:
@@
To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with higher fault detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest [6] to generate a mutation report, as shown in Figure 3. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 4) that guides the LLM in generating more effective test cases during the generation stage. For example, the mutation feedback extracted from the report for the mutant shown in Listing 1 is as follows:
@@ II-B Generation Stage
However, two of the generated test cases fail to execute due to runtime errors. To address this, MutGen incorporates a fixing mechanism that automatically repairs failing test cases, thereby improving both execution success rate and mutation score. In this example, MutGen successfully fixes one failing test case by invoking LLM with the following prompt, resulting in a 13% increase in mutation score.
To further enhance effectiveness, MutGen introduces an iterative generation strategy. In each iteration, additional test cases are generated to target the remaining live and uncovered mutants that were not killed by the previous test suite. After two iterations, for the running example, MutGen achieves a 100% mutation score. During this process, LLMs may occasionally generate test cases with method names that conflict with existing method names. For example, as shown in Listing 2, the LLM generates a test case for the boundary value 04-01-2025 in April. Later, it may generate another test case for 04-31-2025, also intended to test boundary conditions. However, both may be assigned the same method name, leading to naming conflicts. To resolve this, MutGen incorporates an additional guideline, specifically addressing method name conflicts, into the prompt used for fixing (shown in Figure 5).
@@ III-A Overview
Our assumption, as reported in the mutation testing literature [18, 11], is that by maximizing the mutation score, we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion than line or branch coverage [28] and therefore enables the generation of more effective test suites. However, as revealed in recent work [32], finding an optimal prompt for LLMs remains a significant challenge, and we introduce an approximate and iterative solution, detailed in Section III-E, which we investigate empirically and whose limitations we analyze, for example, in terms of mutation operators that are difficult to kill.
@@ III-C Prompting with Mutation Feedback
$\text{Prompt}_{\text{mut}}(\mathcal{R})$ $=\texttt{Format}\left(desc,\bigcup_{r_{i}\in\mathcal{R}}r_{i}\right)$
@@
As illustrated in Figure 6 and the example in Section II-B, the prompt constructed by MutGen includes a natural language code summarization (descdesc), the source code excluding comments (srcsrc), and the associated mutation feedback ($R$). In addition, MutGen explicitly specifies the expected output format from the LLM, as detailed in the prompt example Figure 4 shown in Section II-B.
@@ III-D Fixing Execution Errors
As shown in prior work [72], advanced LLMs are generally capable of producing syntactically correct test cases. However, they often suffer from runtime errors, among which assertion failures are widespread. In this paper, which targets unit testing scenarios, we focus on addressing assertion failures. To this end, MutGen incorporates four fix strategies into the prompt: Direct Invocation, Handling AssertEquals Mismatches, Fixing assertTrue/assertFalse Errors, and Handling Ambiguous Cases. For potential compilation errors, we employ a fifth strategy, which proactively avoids these issues during the process described in Section III-E. For other errors, MutGen adopts a sixth strategy, instructing the LLM to actively repair them rather than ignoring them. All fix strategies are listed in the prompt shown in Figure 5.
@@ III-E Pushing the Limit: Iterative Generation
The process continues until convergence (i.e., no significant improvement in mutation score) or until a predefined iteration limit is reached.
@@ IV Evaluation
As previously discussed, MutGen is designed to enhance the fault-detection capability of LLMs in automated test generation by incorporating mutation feedback. To evaluate its effectiveness, we compare MutGen with three baselines: EvoSuite, the state-of-the-art search-based technique; EvoSuitemut, a variant that replaces coverage-based fitness functions with strong mutation as described in [31]; and Genvanilla, a prompting-based approach that utilizes an LLM without incorporating mutation feedback and leveraging the comments provided in the program under test through summarization (as illustrated in Figure 4). We adopt Llama-3.3 70B as the base model in our evaluation due to its recent release, strong effectiveness on code-related tasks [33], and open-source accessibility, which ensures reproducibility and ease of deployment in both research and practical scenarios.
In addition to effectiveness comparisons, we analyze the limitations of LLM-based test generation by examining killing ratios across different mutation operators. This allows us to identify which types of mutants are particularly challenging for LLMs to kill and offer insights into potential areas for improvement.
@@ IV-A Experiment Setup
Subjects. In our study, we primarily focus on method-level subjects, where each subject is a single Java class containing only one target method. We generate unit test cases specifically for this method. We use a widely adopted benchmark dataset that has been used in prior publications [61, 75, 63, 40], and we also create an additional dataset to enable a more diverse evaluation. More specifically, we include the following subjects:
@@
Following prior work [40, 61, 63, 75], we adopt the HumanEval dataset [19, 14] as one of our evaluation benchmarks. Specifically, we utilize all 160 Java subjects from the dataset. For each subject, we apply EvoSuite with its default configuration to automatically generate a test suite. We then evaluate the mutation score of the generated test suite using PITest [6]. Subjects where MutGen and EvoSuite both achieved a 100% mutation score are discarded, resulting in 104 Java subjects for our evaluation. Indeed, these subjects do not help us distinguish the two approaches.
@@
We search GitHub using the keyword "Leetcode Java" and select the top two repositories based on the Best match sorting criterion. These repositories contain Java solutions to algorithmic problems from LeetCode [3], which are categorized into three difficulty levels: Easy, Medium, and Hard. We exclude all solutions corresponding to Easy problems. We collect all Java source files for the remaining Medium and Hard problems and organize them into a unified package structure following the convention used in HumanEval-Java. We use the Maven build system to compile the project and remove any subjects that failed to compile. For each valid subject, we apply EvoSuite with its default settings to generate a test suite. Subjects for which the generated test suite achieved a 100% mutation score by MutGen and EvoSuite are once again discarded. Finally, for computational feasibility, we randomly select 50 subjects from each of the Medium and Hard sets, resulting in a total of 100 Leetcode-Java subjects used in our evaluation.
@@
Metrics. Since our work focuses on enhancing the effectiveness of LLM-generated test cases, we adopt the mutation score, calculated as the ratio of killed mutants to the total number of mutants, as the primary evaluation metric. A mutant is considered killed if at least one generated test case causes a behavioral difference between the original program and its mutated version, typically resulting in a failed assertion or exception. In addition to the mutation score, we also report branch coverage and line coverage, in line with previous work [8, 20].
@@
Process. To address RQ1, we use all subjects from both datasets. We begin by generating test cases with both Genvanilla and MutGen through a single end-to-end execution, recording the execution time and the number of tokens consumed. We then apply EvoSuite and EvoSuitemut to generate test cases for each subject, setting their time budget to the larger average processing times of MutGen on the two datasets. It is important to note that, unlike prior work [28], our approach generates tests entirely from scratch, without relying on any existing test cases. We evaluate each test suite using two structural coverage metrics, branch coverage and line coverage, measured by the JaCoCo tool [2]. Additionally, we use PITest [6] with its DEFAULTS mutation operator group to measure the mutation score of each test suite. While PITest also offers a STRONGER operator group, it differs from DEFAULTS by only two additional mutation operators, whose applicability is rather limited. For example, the Remove Conditionals operator only targets equality checks. As for the ALL group provided by PITest, the extra mutation operators in this group generally generate mutants that are relatively easy to kill, making them less informative for our evaluation. To assess the statistical significance of the improvements achieved by our approach in terms of effectiveness, we performed non-parametric hypothesis testing using Vargha and Delaney’s A12 statistic [13, 64]. The A12 effect size was computed based on the mutation score, which ranges from 0 to 1: a value of 0.5 indicates no difference (i.e., results are due to chance), while a value greater than 0.5 suggests that MutGen is more likely to outperform the compared approach. Conversely, a value less than 0.5 indicates the opposite. We also conducted a paired-sample Wilcoxon signed-rank test on both coverage and mutation score to evaluate whether MutGen achieves statistically significant improvements on these metrics.
@@
To address RQ2, we analyze the mutation operators involved in the evaluation. As a preliminary experiment, we first randomly select 10 subjects from each dataset and perform seven rounds of end-to-end test generation. In each round, MutGen generates additional test cases based on the previously generated ones. By tracking the mutation scores across iterations, we observe that the scores tend to converge after the fourth iteration. Therefore, we set the maximum number of repetitions to four in our experiments. We collect results at the fourth iteration, and analyze the distribution of mutation operators across all generated mutants and compute the killing ratio for each operator, that is, the proportion of killed mutants relative to the total number of mutants produced by that operator. This allows us to identify which mutation operators tend to produce mutants that are difficult or easy for LLM-generated test cases to kill. In our evaluation, mutation testing is conducted using PITest, which employs 11 mutation operators as described in [1]. For mutation operators with low killing ratios, we investigate why LLMs struggle to kill these mutants. Due to space constraints, we do not present a detailed analysis for each mutation operator or for all live and uncovered mutants. Instead, we analyze a subset of cases and identify several factors contributing to the inability of LLMs to kill certain mutants. Nevertheless, our findings offer valuable insights and highlight important directions for future research in this area.
@@
Implementation. We describe important implementation details of MutGen and our evaluation setup. All subjects used in the evaluation are built using the Maven build system. We follow official guidelines to integrate EvoSuite, JaCoCo, and PITest into the Maven pipeline for automated test generation, code coverage measurement, and mutation testing, respectively. Our approach, MutGen, is implemented in Python and built on top of the Mutahunter framework [4]. Specifically, we reuse components from Mutahunter for prompt handling, LLM interfacing, and response processing. We use Ollama [5] to locally deploy the Llama-3.3 70B model [33] with the default configuration.

### v5

@@ I Introduction
Unit test generation is a fundamental task in software engineering, crucial for ensuring code quality and reducing manual testing effort [naik2011software, fraser2014large, corno2004automatic, anand2013orchestrated]. With the rapid advancement of Large Language Models (LLMs), research has shifted from traditional techniques—such as random testing [pacheco2007randoop, hamlet1994random, andrews2011genetic, oriat2005jartege, pacheco2007feedback] and search-based methods [fraser2011evosuite, rojas2016seeding, rojas2015combining, vivanti2013search]—to LLM-based approaches for generating unit tests [fan2023large, wang2024software, hou2024large].
@@
However, most existing work evaluates generated test cases primarily based on code coverage metrics, such as line or branch coverage. According to prior studies [chekam2017empirical, andrews2006using, inozemtseva2014coverage], high code coverage does not necessarily imply strong fault detection capability. Other research [chekam2017empirical, alshahwan2024automated, foster2025mutation] suggests that the mutation score is a more reliable and meaningful metric for evaluating the effectiveness of test cases. Mutation testing introduces artificial faults, known as mutants, into the program by applying syntactic changes using mutation operators. The generated test cases are executed against the mutants, and if a test causes a mutant to fail (i.e., “kills” it), it is considered effective. The mutation score is the ratio of killed mutants to the total number of generated mutants.
@@
Recent efforts have begun exploring mutation-based evaluation for LLM-generated test cases. For example, Foster et al. [foster2025mutation] investigate regression testing scenarios in privacy-critical Kotlin applications such as WhatsApp and Instagram. They focus on generating test cases from existing ones and generating mutants using LLMs, which introduces challenges such as detecting equivalent mutants that no test can kill. Another work by Dakhel et al. [dakhel2024effective] targets Python benchmarks and proposes a prompt-based approach to improve mutation score. However, they iteratively add a single mutant until they cannot kill the new mutant. Therefore, they do not attempt to converge towards a maximum mutation score to maximize fault detection. Moreover, their study does not analyze why live mutants remain undetected, nor does it consider uncovered mutants, both of which are challenges we address in our research.
@@
While these efforts represent promising first steps, the use of mutation testing as a core evaluation and improvement strategy for LLM-generated test cases remains underexplored. To address this gap, we propose MutGen, a mutation-guided LLM-based approach that explicitly maximizes the mutation score. Specifically, a key feature of MutGen is that we incorporate mutation feedback (i.e., live and uncovered mutant information) into prompt construction to guide the LLM in generating more effective test cases. Further, as discussed in prior studies [tian2025fixing, wang2025projecttest, pan2024multi], incorporating a fixing step is crucial for improving test generation effectiveness. While prior work primarily focuses on improving metrics such as code coverage or test passing rate, our work explicitly targets the fault detection capability of test cases, measured by the mutation score. To this end, our approach introduces a fixing step that specifically addresses assertion failures in the generated test cases, caused by incorrect use of assertion functions, including erroneous invocations and inappropriate parameter values.
@@
Moreover, we observe that comments in the target code can mislead LLMs into producing irrelevant or incorrect tests. To address this, we introduce a code summarization step during preprocessing that replaces misleading comments with concise summaries, which are then embedded into the prompt. Lastly, since achieving the maximum mutation score is inherently difficult, we introduce an iterative test-generation strategy to push LLMs toward killing more live mutants. Once the mutation score converges, we analyze both uncovered and live mutants, as well as specific mutation operators that remain particularly challenging for the LLM. Like many recent studies [yuan2024evaluating, gu2025llm, ryan2024code], our approach focuses on function-level testing. While this prior work also generates test cases for functions from complex datasets such as Defects4J [just2014defects4j], it handles cross-module dependencies explicitly by simply providing them in the prompts. In contrast, our work does not aim to address dependency resolution at this stage; instead, it focuses on improving mutation scores for function-level testing. Nevertheless, as discussed below, one of our datasets includes solutions to challenging algorithmic problems with complex program structures, providing a more complex dataset for evaluating the effectiveness of MutGen in generating test cases.
@@
To evaluate MutGen, we use two datasets: HumanEval-Java, widely adopted in prior work [tian2025fixing, dakhel2024effective, siddiq2024using], and a new LeetCode-Java corpus containing 100 randomly selected algorithmic problems with ground-truth solutions. The experimental results show that test cases generated by MutGen outperform both EvoSuite, a state-of-the-art search-based technique, and a vanilla LLM-based method in terms of mutation score. MutGen achieves mutation scores of 89.5% and 89.1% on HumanEval-Java and LeetCode-Java, respectively, across a total of 1,144 and 1,900 mutants.
@@ II Motivating Example and Informal Description
When combined with the example code (including comments), LLMs can generate test cases that achieve high line and branch coverage. However, as demonstrated in prior work [foster2025mutation, chekam2017empirical, alshahwan2024automated], high coverage does not necessarily imply strong fault detection capability, for example, when measured as mutation score. For instance, in our experiments, LLMs generate tests for the subject id_81 from HumanEval-Java with 100% line and branch coverage, yet the corresponding mutation score is only 4%.
Listing 1 shows a representative test function generated by Llama-3.3. In practice, multiple such test functions may be generated, often with a variety of data inputs aimed at maximizing code coverage. However, we observe that most of these inputs fail to cover corner cases such as February 29th on leap years, likely due to the absence of such examples in the provided comment block. The comments in the target code may also introduce additional issues. For example, the comment begins with “You have to write a function…”, which, when included as part of the prompt, may mislead the LLM into replicating the target code logic rather than generating a meaningful test function. Additionally, Llama-3.3 occasionally produces dates in incorrect formats such as dd-mm-yyyy, likely due to ambiguity or insufficient specification of the expected format in the comments provided within the program under test.
Further, when relying on mutation instead of code coverage, the generated inputs often fail to kill mutants effectively. For example, when the ConditionalsBoundary operator in PITest changes a conditional boundary and alters line 24 to:
@@
Listing 1: Example test function generated by Llama-3.3
@@ II-A Preprocessing Stage
For the target code shown in Figure 1, Llama-3.3 returns the following summary:
@@
To address the second issue, we leverage mutation feedback to guide the LLM in generating test cases with higher fault detection capability. Before the generation stage, MutGen extracts mutation feedback from mutation reports produced by mutation testing tools. In this example, we use PITest [pitest] to generate a mutation report, as shown in Figure 3. From the report, MutGen retrieves key information, including the mutation location (line number), the mutant status (killed, survived, or uncovered), and the mutation operator applied to create the mutant. MutGen then integrates this feedback with the source code, forming a prompt (shown in Figure 4) that guides the LLM in generating more effective test cases during the generation stage. For example, the mutation feedback extracted from the report for the mutant shown in Listing II is as follows:
@@ II-B Generation Stage
However, two of the generated test cases fail to execute due to runtime errors. To address this, MutGen incorporates a fixing mechanism that automatically repairs failing test cases, thereby improving both execution success rate and mutation score. In this example, MutGen successfully fixes one failing test case by invoking LLM with the prompt shown in Figure 5, resulting in a 13% increase in mutation score.
To further enhance effectiveness, MutGen introduces an iterative generation strategy. In each iteration, additional test cases are generated to target the remaining live and uncovered mutants that were not killed by the previous test suite. After two iterations, for the running example, MutGen achieves a 100% mutation score. During this process, LLMs may occasionally generate test cases with method names that conflict with existing method names. For example, as shown in Listing 1, the LLM generates a test case for the boundary value 04-01-2025 in April. Later, it may generate another test case for 04-31-2025, also intended to test boundary conditions. However, both may be assigned the same method name, leading to naming conflicts. To resolve this, MutGen incorporates an additional guideline, specifically addressing method name conflicts, into the prompt used for fixing (shown in Figure 5).
@@ III-A Overview
Our assumption, as reported in the mutation testing literature [chekam2017empirical, andrews2006using], is that by maximizing the mutation score, we also maximize fault detection effectiveness. In any case, achieving mutation coverage is known to be a more stringent criterion than line or branch coverage [foster2025mutation] and therefore enables the generation of more effective test suites. However, as revealed in recent work [gao2025prompt], finding an optimal prompt for LLMs remains a significant challenge, and we introduce an approximate and iterative solution, detailed in Section III-E, which we investigate empirically and whose limitations we analyze, for example, in terms of mutation operators that are difficult to kill.
@@ III-B Preprocessing
At the start of our approach, all mutants are initially alive, and the items of mutation feedback are formatted as shown in the above example.
@@ III-C Prompting with Mutation Feedback
$\text{Prompt}_{\text{mut}}(\mathcal{R})$ $=\texttt{Format}\left(desc,{src,}\bigcup_{r_{i}\in\mathcal{R}}r_{i}\right)$
@@
As illustrated in Figure 6 and the example in Section II-B, the prompt constructed by MutGen includes a natural language code summarization (descdesc), the source code excluding comments (srcsrc), and the associated mutation feedback ($R$). In Figure 4, these correspond to the values of the prompt variables: descdesc is represented as {{summary}}, srcsrc as {{source code without comments}}, and $\mathcal{R}$ as the content under ## Mutation Feedback. In addition, MutGen explicitly specifies the expected output format from the LLM, as detailed in the prompt example Figure 4 shown in Section II-B.
@@ III-D Fixing Execution Errors
As shown in prior work [yuan2024evaluating], advanced LLMs are generally capable of producing syntactically correct test cases. However, they often suffer from runtime errors, among which assertion failures are widespread. In this paper, which targets unit testing scenarios, we focus on addressing assertion failures. To this end, MutGen incorporates four fix strategies into the prompt: Direct Invocation, Handling AssertEquals Mismatches, Fixing assertTrue/assertFalse Errors, and Handling Ambiguous Cases. For potential compilation errors, we employ a fifth strategy, which proactively avoids these issues during the process described in Section III-E. For other errors, MutGen adopts a sixth strategy, instructing the LLM to actively repair them rather than ignoring them. All fix strategies are listed in the prompt shown in Figure 5.
@@ III-E Pushing the Limit: Iterative Generation
The process continues until convergence (i.e., no significant improvement in mutation score) or until a predefined iteration limit is reached. Once convergence is reached, the remaining live and uncovered mutants provide insights into mutation operators and program constructs that remain challenging for the LLM, which we analyze in Section IV-B.
@@ IV Evaluation
As previously discussed, MutGen is designed to enhance the fault-detection capability of LLMs in automated test generation by incorporating mutation feedback. To evaluate its effectiveness, we compare MutGen with three baselines: EvoSuite, the state-of-the-art search-based technique; EvoSuitemut, a variant that replaces coverage-based fitness functions with strong mutation as described in [fraser2015achieving]; and Genvanilla, a prompting-based approach that utilizes an LLM without incorporating mutation feedback and leveraging the comments provided in the program under test through summarization (as illustrated in Figure 4). We adopt Llama-3.3 70B as the base model in our evaluation due to its recent release, strong effectiveness on code-related tasks [grattafiori2024llama], and open-source accessibility, which ensures reproducibility and ease of deployment in both research and practical scenarios.
In addition to effectiveness comparisons, we analyze the limitations of LLM-based test generation by examining killing ratios across different mutation operators. This allows us to identify which mutant types are particularly challenging for LLMs to kill and to offer insights into potential areas for improvement.
@@ IV-A Experiment Setup
Subjects. In our study, we primarily focus on method-level subjects, where each subject is a single Java class containing only one target method. We generate unit test cases specifically for this method. We use a widely adopted benchmark dataset that has been used in prior publications [siddiq2024using, zheng2023codegeex, tian2025fixing, islam2405mapcoder], and we also create an additional dataset to enable a more diverse evaluation. More specifically, we include the following subjects:
@@
Following prior work [islam2405mapcoder, siddiq2024using, tian2025fixing, zheng2023codegeex], we adopt the HumanEval dataset [chen2021evaluating, athiwaratkun2022multi] as one of our evaluation benchmarks. Specifically, we utilize all 160 Java subjects from the dataset. For each subject, we apply EvoSuite with its default configuration to automatically generate a test suite. We then evaluate the mutation score of the generated test suite using PITest [pitest]. Subjects where MutGen and EvoSuite both achieved a 100% mutation score are discarded, resulting in 104 Java subjects for our evaluation. Indeed, these subjects do not help us distinguish the two approaches.
@@
We search GitHub using the keyword "Leetcode Java" and select the top two repositories based on the Best match sorting criterion. These repositories contain Java solutions to algorithmic problems from LeetCode [leetcode], which are categorized into three difficulty levels: Easy, Medium, and Hard. We exclude all solutions corresponding to Easy problems. We collect all Java source files for the remaining Medium and Hard problems and organize them into a unified package structure following the convention used in HumanEval-Java. We use the Maven build system to compile the project and remove any subjects that failed to compile. For each valid subject, we apply EvoSuite with its default settings to generate a test suite. Subjects for which the generated test suite achieved a 100% mutation score by MutGen and EvoSuite are once again discarded. Finally, for computational feasibility, we randomly select 50 subjects from each of the Medium and Hard sets, resulting in a total of 100 Leetcode-Java subjects used in our evaluation.
The average number of lines of code in HumanEval-Java methods is 41, with an average cyclomatic complexity (CC) of 4.90, which is close to the average CC of 5.46 reported for typical Java methods [cc]. This indicates that HumanEval-Java exhibits representative control-flow complexity for standard Java methods. In contrast, methods in Leetcode-Java are generally more complex, with an average CC of 7.88. These statistics demonstrate that Leetcode-Java contains methods with complexity levels comparable to or exceeding those found in real-world Java code [ccpaper], making it both representative and more challenging than HumanEval-Java.
@@
Metrics. Since our work focuses on enhancing the effectiveness of LLM-generated test cases, we adopt the mutation score, calculated as the ratio of killed mutants to the total number of mutants, as the primary evaluation metric (see Equation 1).
$\text{Mutation Score}=\frac{\text{\# killed mutants}}{\text{\# total mutants}}\times 100\%$ (1)
A mutant is considered killed if at least one generated test case causes a behavioral difference between the original program and its mutated version, typically resulting in a failed assertion or exception. In addition to the mutation score, we also report branch coverage and line coverage, in line with previous work [alagarsamy2024a3test, chen2024chatunitest].
@@
Process. To address RQ1, we use all subjects from both datasets. We begin by generating test cases with both Genvanilla and MutGen through a single end-to-end execution, recording the execution time and the number of tokens consumed. We then apply EvoSuite and EvoSuitemut to generate test cases for each subject, setting their time budget to the larger average processing times of MutGen on the two datasets. It is important to note that, unlike prior work [foster2025mutation], our approach generates tests entirely from scratch, without relying on any existing test cases. We evaluate each test suite using two structural coverage metrics, branch coverage and line coverage, measured by the JaCoCo tool [jacoco]. Additionally, we use PITest [pitest] with its DEFAULTS mutation operator group to measure the mutation score of each test suite. The DEFAULTS group provides a set of operators commonly used in the literature [fraser2015achieving, wang2024comprehensive] and, according to the PITest official documentation [pitest], are not easy to detect and minimize the number of equivalent mutants. While PITest also offers a STRONGER operator group, it differs from DEFAULTS by only two additional mutation operators, whose applicability is rather limited. For example, the Remove Conditionals operator only targets equality checks. As for the ALL group provided by PITest, the extra mutation operators in this group generally generate mutants that are relatively easy to kill, making them less informative for our evaluation. Note that although PITest is designed for Java, its mutation operators can be applied to other languages as well [pitest, wang2024comprehensive, jia2010analysis]. To assess the statistical significance of the improvements achieved by our approach in terms of effectiveness, we performed non-parametric hypothesis testing using Vargha and Delaney’s A12 statistic [arcuri2011practical, vargha2000critique]. The A12 effect size was computed based on the mutation score, which ranges from 0 to 1: a value of 0.5 indicates no difference (i.e., results are due to chance), while a value greater than 0.5 suggests that MutGen is more likely to outperform the compared approach. Conversely, a value less than 0.5 indicates the opposite. We also conducted a paired-sample Wilcoxon signed-rank test on both coverage and mutation score to evaluate whether MutGen achieves statistically significant improvements on these metrics.
@@
To address RQ2, we analyze the mutation operators involved in the evaluation. As a preliminary experiment, we first randomly select 10 subjects from each dataset and perform seven rounds of end-to-end test generation. In each round, MutGen generates additional test cases from previously generated ones. By tracking mutation scores across iterations, we observe that they tend to converge after the fourth iteration. Therefore, we set the maximum number of repetitions to four in our experiments. We collect results at the fourth iteration, and analyze the distribution of mutation operators across all generated mutants and compute the killing ratio for each operator, that is, the proportion of killed mutants relative to the total number of mutants produced by that operator. This allows us to identify which mutation operators tend to produce mutants that are difficult or easy for LLM-generated test cases to kill. In our evaluation, mutation testing is conducted using PITest, which employs 11 mutation operators as described in [mutops]. For mutation operators with low killing ratios, we investigate why LLMs struggle to kill these mutants. Due to space constraints, we do not present a detailed analysis for each mutation operator or for all live and uncovered mutants. Instead, we analyze a subset of cases and identify several factors that contribute to LLMs’ inability to kill certain mutants. Nevertheless, our findings offer valuable insights and highlight important directions for future research in this area.
@@
Implementation. We describe important implementation details of MutGen and our evaluation setup. All subjects used in the evaluation are built using the Maven build system. We follow official guidelines to integrate EvoSuite, JaCoCo, and PITest into the Maven pipeline for automated test generation, code coverage measurement, and mutation testing, respectively. Our approach, MutGen, is implemented in Python and built on top of the Mutahunter framework [mutahunter]. Specifically, we reuse components from Mutahunter for prompt handling, LLM interfacing, and response processing. We use Ollama [ollama] to locally deploy the Llama-3.3 70B model [grattafiori2024llama] with the default configuration, setting the temperature to 0.0 for more deterministic outputs. As our main baseline, EvoSuite officially and stably supports JUnit 4, while its support for JUnit 5 is currently limited and experimental. Therefore, JUnit 4 remains the recommended choice for reliable test generation. To ensure a fair comparison, we adopt JUnit 4 in all experimental settings.